If 2 * 4 = 10, 3 * 5 = 13, what is 4 * 7?
Please help me with solution pleaseeee;

Found 2 solutions by josgarithmetic, addingup:
Answer by josgarithmetic(39799)

You may need to try to figure what the * asterisk symbol means. The way to find something that works will need effort to look for anything possible that is consistant.

Multiplying or adding did not lead to anything too clear.

2*4=10?
(2+4)/2=3, the average; (3)(3)=9; 9+1=10.

3*5=13?
(3+5)/2=4, (4)(3)=12; 12+1=13.

The way your asterisk operation works may be for any two numbers, x and y,
x*y=((x+y)/2)(3)+1.
_
According to this as the rule for x*y,
4*7=((4+7)/2)(3)+1=(11/2)(3)+1=33/2+1=33/2+2/2=35/2

The answer is probably 4*7=33/2
according to the rule for the asterisk operation found from your two examples.

Answer by addingup(3677)

what did they tell you to do?
2*4=8+2=10
3*5=15-2=13
4*7=28+?or-? If you heard that patterns repeat, then the next number is 28+2=30 but you were in class, I wasn't, so I don't know what you were told.